# Pontryagin Optimal Control via Neural Networks

### 07 Jan 2024 v2

@@ IV Numerical Examples
In this section, we evaluate the optimal control performance of our proposed NN-PMP-Gradient controller and illustrate its versatile capabilities on numerical simulations of 5 tasks. In simulations we benchmark against: i). a Linear-Convex solver, which we firstly use linear regression to approximate a linearized system dynamics and then solve optimal control problem using cvxpy solver [28]; ii). state-of-the-art model-free RL algorithm Proximal Policy Optimization (PPO); and iii). model-based random shooting MPC (RS-MPC) controller (based on NN-simulated system dynamics) [17]. To benchmark performances, for all simulations we use a Google Colab T4 GPU with 12GB RAM.
IV-A Experiment Control Tasks
In this subsection, we briefly introduce 5 optimal control environments for our experiment. The settings of each task (including time horizon $T$, state and action dimension $m,n$, running cost $L(\mathbf{x}_{t},\mathbf{u}_{t})$, terminal cost $\Phi(\mathbf{x}_{T})$ and system dynamics $f(\mathbf{x}_{t},\mathbf{u}_{t})$ is illustrated in Table. I. We note that in all test cases, it is assumed that the ground truth dynamics are not revealed to the controller. Detailed settings of these systems can be found in Appendix C in our paper’s online version [27].
Figure 2: Sample-Performance (Cost/Reward) curves on 5 different experiment environments. (a) LQR (Cost); (b) Battery (Cost); (c) Pendulum (Cost); (d) MuJoCo Swimmer (Reward); (e) MuJoCo HalfCheetah (Reward).
2. Battery. We consider optimal arbitrage of a battery. The battery make profits via participating in electricity markets and making charging or discharging decisions under time-varying prices. A profitable energy storage arbitrage plan is crucial, since it encourages utilization of stochastic renewables and addressing energy sustainability [29]. The battery has a nonlinear and unknown charging efficiency with respect to battery’s state-of-charge.
IV-B Numerical Results
We evaluate the performance of our proposed NN-PMP-Gradient on the 5 tasks mentioned in Section IV-A. For LQR, Single Pendulum, Swimmer and HalfCheetah experiments, we apply neural network $f_{NN}$ to learn the full system dynamics $f(\mathbf{x}_{t},\mathbf{u}_{t})$, while in Battery case we use $f_{NN}$ to learn part of the dynamics: the charging efficiency function $\zeta(\mathbf{u}_{t})$. In all 5 experiments, we compare our NN-PMP-Gradient method against Linearized model, classic model-free RL method PPO and model-based RL method RS-MPC (based on the same NN-simulated system dynamics in our NN-PMP-Gradient method). For all environments tested we find NN-PMP-Gradient converges during policy inference.
LQR
Battery
Pendulum
Swimmer
HalfCheetah
@@ IV-B Numerical Results
TABLE II: Performances under varying training samples, according to the complexity of experiment environments. The performances are averaged for 10 runs test. For LQR, Battery and Pendulum the control performance is evaluated by control costs (the lower the better); for Swimmer and HalfCheetah we evaluate the accumulated reward (the higher the better).
Table. II shows the control performance of each control method across all 5 experiments. In general, the results demonstrate that our NN-PMP-Gradient outperforms Linearized model, PPO, and RS-MPC in terms of control performance using same number of training samples. For example, in Single Pendulum experiment, with NN-dynamics trained by 2,000 samples, our NN-PMP-Gradient reduces the cost by at least 90 compared with Linearized model, PPO and RS-MPC, yielding more than 8.2% improvement; while in HalfCheetah case, with NN-dynamics trained by 100,000 samples, our method shows more than 200 leading in reward, which is huge improvement compared to the best result $96.73$ (achieved by PPO) among the rest 3 methods. We also note that in specific cases with strong local optima like Swimmer [31], though NN-PMP-Gradient still achieves competitive control performance, lack of random search may add difficulty for escaping from local optima, making it underperform model-free PPO with a very narrow gap. Meanwhile, the inaccuracy when using NN to approximate complex, high-dimension system dynamics like HalfCheetah, causes bigger variances in NN-PMP-Gradient’s control performance. We regard such observations as limitations of our method, and will address them in the future work.
Fig. 2 illustrates the control performance trend of each method as the training samples increase. We find our NN-PMP-Gradient keeps a considerable leading in control performance. In most cases, even with significantly fewer training samples for learning the system dynamics, our method still achieves comparable results with higher reward/lower costs.
@@
TABLE III: State Bounds Exceeding Rates (SBER) on the Battery experiment. Results are averaged for 3 runs (each contains 10 trials).
@@
Figure 3: A State Trajectory Example in Battery experiment. Only trajectory obtained by NN-PMP-Gradient does not exceed state bounds, while others does.
@@
Furthermore, simulation results highlight advantages of our NN-PMP-Gradient on satisfying safety constraints for control problems defined by the state or action constraints. In the Battery case, Table. III shows the average State Bounds Exceeding Rate (SBER) in battery experiment with 2,000 training samples, which records the proportion of state trajectories that exceed bounds obtained by various controllers across total time steps. Fig. 3 shows one-day state trajectory comparison for the Battery case. We can observe only our method constrains the battery state-of-charge within the pre-defined $0-10$ kWh limits, while other algorithms’ resulting actions cause the battery to be easily overdischarged with state-of-charge smaller than 00 kWh. More importantly, we can observe our proposed algorithm can charge the battery when price is lower, while selling power at hours with higher price during noon. These results further indicate that our NN-PMP-Gradient yields safer and realistic solutions, and hence improves the control performance significantly.
@@ V Conclusion and Discussion
In this paper, we showed that the classic theories of optimal control, the Pontryagin’s Principle, can be seamlessly integrated into data-driven settings. The proposed NN-PMP-Gradient is able to both accurately simulate and optimally control unknown dynamical systems. Simulations on five numerical examples (LQR, battery, single pendulum, MuJoCo Swimmer and HalfCheetah) demonstrated our NN-PMP-Gradient outperforms standard optimization-based solver using linearized dynamics, model-free PPO and model-based RS-MPC under a variety of settings. For complex real-world systems, the proposed method is a more sample-efficient, generalizable and safer optimal controller.
There exist several limitations to the current PMP-based data-driven control scheme. The current method can be affected by modeling errors during NN training, which causes high variances in control performance for high-dimensional systems like MuJoCo HalfCheetah. Additionally, for some specific scenarios with strong local optima like Swimmer, lack of random search limits proposed method’s performance. Moreover, for problems with long horizon and high dimension, the co-state recursion in PMP-Gradient can impose high computational burden. In the future, we aim to address these challenges and seek theoretical insights to enhance our NN-PMP-Gradient as a more robust and efficient model-based optimal controller.

### 15 Jan 2024 v3

@@ IV Numerical Examples
LQR (C)
Battery (C)
Pendulum (C)
Swimmer (R)
HalfCheetah (R)
@@
TABLE II: Performances under varying number of training samples. The performances are averaged for 10 runs test.
@@
TABLE III: State Bounds Exceeding Rates (SBER) on the Battery simulation. Results are averaged for 3 runs (each contains 10 trials).
In this section, we evaluate the optimal control performance of our proposed Neural-PMP controller and illustrate its versatile capabilities on numerical simulations of 5 tasks. In simulations we benchmark against: i). a Linear-Convex solver, which we firstly use linear regression to approximate a linearized system dynamics and then solve MPC problem using cvxpy solver; ii). state-of-the-art model-free RL algorithm Proximal Policy Optimization (PPO); and iii). model-based random shooting MPC (RS-MPC) controller (based on NN-simulated system dynamics) [16]. To benchmark performances, for all simulations we use a Google Colab T4 GPU with 12GB RAM.
IV-A Simulated Control Tasks
The settings of 5 benchmarked control tasks (including time horizon $T$, state and action dimension $m,n$, running cost $L(\mathbf{x}_{t},\mathbf{u}_{t})$, terminal cost $\Phi(\mathbf{x}_{T})$ and system dynamics $f(\mathbf{x}_{t},\mathbf{u}_{t})$ are illustrated in Table. I. We note that in all test cases, it is assumed that the ground truth dynamics are not revealed to the controller. Detailed settings of these systems can be found in Appendix C in our paper’s online version [27]. Our codes for the controller and numerical simulations are publicly available at https://github.com/ChengyangGU/NeuralPMP2024.
Figure 2: Sample-Performance (Cost/Reward) curves on (a) LQR (C); (b) Battery (C); (c) Pendulum (C); (d) Swimmer (R); (e) HalfCheetah (R).
2. Battery. We consider optimal arbitrage of a battery. The battery maximizes profits via participating in electricity markets and making charging or discharging decisions under time-varying prices. A profitable energy storage arbitrage plan is crucial, since it encourages utilization of stochastic renewables and addressing energy sustainability [28]. The battery has a nonlinear and unknown charging efficiency with respect to battery’s state-of-charge.
IV-B Numerical Results
We evaluate the control performance, sample-efficiency and safety of our proposed Neural-PMP on the 5 tasks mentioned in Section IV-A. For LQR, Single Pendulum, Swimmer and HalfCheetah simulations, we apply neural network $f_{NN}$ to learn the full system dynamics $f(\mathbf{x}_{t},\mathbf{u}_{t})$, while in Battery case we use $f_{NN}$ to learn part of the dynamics: the charging efficiency function $\zeta(\mathbf{u}_{t})$. In all 5 simulations, we compare our Neural-PMP method against Linearized model, classic model-free RL method PPO and model-based RL method RS-MPC (based on the same NN-simulated system dynamics in our Neural-PMP method). For all environments tested we find Neural-PMP converges during policy inference.
Table. II shows the control performance of each control method across all 5 environments. In general, the results demonstrate that our Neural-PMP outperforms Linearized model, PPO, and RS-MPC in terms of control performance using same number of training samples. For example, in Single Pendulum simulation, with NN-dynamics trained by 2,000 samples, our Neural-PMP reduces the cost by at least 90 compared with Linearized model, PPO and RS-MPC, yielding more than 8.2% improvement; while in HalfCheetah case, with NN-dynamics trained by 100,000 samples, our method shows more than 200 leading in reward, which is huge improvement compared to the best result $96.73$ (achieved by PPO) among the rest 3 methods. We also note that in specific cases with strong local optima like Swimmer [30], though Neural-PMP achieves competitive control performance, it can converge to local optima from some initialized actions, making it underperform model-free PPO with a very narrow gap. We will further address these concerns in the future work.
@@ IV-B Numerical Results
Figure 3: Battery state trajectory obtained by Neural-PMP is the only one not exceeding state bounds.
@@
To validate the sample efficiency, Fig. 2 demonstrates that our method maintains a significant advantage in control performance with varying number of training data. In most cases, even with substantially fewer training samples for learning the system dynamics, our method achieves higher rewards/ lower costs. Furthermore, simulation results highlight advantages of our Neural-PMP on satisfying safety constraints defined by the state or action constraints. Table. III shows the average State Bounds Exceeding Rate (SBER) in battery simulation with 2,000 training samples, which records the proportion of state trajectories that exceed bounds obtained by various controllers across total time steps. Fig. 3 shows one-day state trajectory comparison for the Battery case. We can observe only our method constrains the battery state-of-charge within the pre-defined $0-10$ kWh limits, while other algorithms’ resulting actions cause the battery to be easily overdischarged with state-of-charge smaller than 00 kWh. More importantly, unlike the irregular solution obtained by PPO or RS-MPC, our proposed algorithm can charge the battery when price is lower, while selling power at hours with higher price during noon, which is more interpretable. These results further indicate that our Neural-PMP yields safer and realistic solutions, and hence improves the control performance significantly. In summary, our Neural-PMP outperforms other benchmarked methods in terms of control performance and sample efficiency across all 5 environments, and can achieve safer solutions. These advantages of our Neural-PMP is attributed to its ability to learn the unrevealed dynamics while utilizing PMP conditions for finding optimal control sequences.
@@ V Conclusion and Discussion
In this paper, we showed that the classic Pontryagin’s Principle can be seamlessly integrated into data-driven settings. The proposed Neural-PMP is able to both accurately simulate and optimally control unknown dynamical systems. Simulations on five numerical examples demonstrated our Neural-PMP outperforms standard optimization-based solver using linearized dynamics, model-free and model-based RL under a variety of settings. For complex real-world systems, the proposed method is a more sample-efficient, generalizable and safer optimal controller. For future work, we will investigate the effects of NN modeling errors and initialization of policies in our Neural-PMP controller design. We aim to address these challenges and seek theoretical insights to enhance our Neural-PMP as a more robust and efficient model-based optimal controller.